OpenStudy (anonymous):

quadratic equations : y(y+1) = 20?

OpenStudy (anonymous):

y^2 + y = 20 y^2 + y - 20 = 0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

factor y^2 + y - 20 = 0

OpenStudy (anonymous):

No idea

OpenStudy (anonymous):

-20 = 5 * -4 5 - 4 = 1

OpenStudy (turingtest):

you mean y^2 + y - 20 =(y+5)(y-4)

OpenStudy (turingtest):

and let's stop here, if the questioner can't move past this we shouldn't give the answer

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

-5 4

OpenStudy (turingtest):

yes! but try to write this as x={-5,4} that is the proper notation

OpenStudy (turingtest):

try to follow what moneybird did, you will be getting weened off our help slowly and will be expected to answer more on your own in the future